# Solitons in binary compounds with stacked two-dimensional honeycomb lattices

### v1

@@ IV.1.3 Soliton-antisoliton pair
Figure 9: Low-energy band structures for ${}^{\prime}$ stacking. Each column shows a different value of the onsite energies $U$. The top row shows a faultless system with $N=16$ layers, whereas the second row shows $N=16$ layers with a sharp fault at its center $(8,8)$. In all plots, parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [35], $a=2.46$Å [36].
IV.2 ${}^{\prime}$
@@ IV.2.1 Faultless system
For bulk, faultless ${}^{\prime}$ stacking, using the approximation for the intralayer hopping (23) with $k<k_{c}$, we can use a unit cell of two atoms, either $A_{i}$, $B_{i}$ on an odd layer or $B_{i}$, $A_{i}$ on an even layer, to give the Hamiltonian
@@ IV.2.2 Single soliton
Band structures for ${}^{\prime}$ stacking with a sharp fault at its center are shown in Fig. 9 (bottom row). At $k=0$, the system separates into isolated atoms with energies ±$\pm U$ plus the $N$-mer chain. Now the $N$-mer chain contains a fault, so it consists of a monoatomic chain with onsite energies $-U$ and nearest neighbor hopping $\gamma_{1}$ connected to a monoatomic chain with onsite energies $+U$ and nearest neighbor hopping $\gamma_{1}$. The former has a bulk band $E(q_{z})=-U+2\gamma_{1}\cos(q_{z}d)$, the latter has a bulk band $E(q_{z})=U+2\gamma_{1}\cos(q_{z}d)$, i.e., the system consists of one part with a band centered on energy $+U$ and of bandwidth $4\gamma_{1}$, connected to another part with a band centered on energy $-U$ and of bandwidth $4\gamma_{1}$. For $U\leq 2\gamma_{1}$, the two bands overlap leading, in a system of a finite number of layers, to overlapping bands with anticrossings at low energy as in Fig. 9(h) to Fig. 9(j). For $U>2\gamma_{1}$, the two bands do not overlap, giving an overall band gap as in Fig. 9(l).
Although the presence of a fault for ${}^{\prime}$ stacking creates changes in the band structure, Fig. 9, there are no states localized on the fault. This can be understood by considering the system at $k=0$: The fault divides two sections of monoatomic chains with different onsite energies ($+U$ or $-U$) and, thus, the two sections are not degenerate, nor do they have a Dirac-like low-energy continuum description [25].
@@ V.1.2 Single soliton
As with ${}^{\prime}$ stacking, although the presence of a fault for AA stacking creates changes in the band structure, Fig. 10, there are no states localized on the fault. The term describing hopping along the ladder, $2\gamma_{1}\cos(q_{z}d)$, appears on the main diagonal in the ladder Hamiltonian (114) and it breaks chiral symmetry. The trivial nature of the fault may be further understood by considering the limit $k=0$ when the system separates into monoatomic chains which do not have a Dirac-like low-energy continuum description [25].
Figure 11: ${}^{\prime}$ stacking with $N=16$ layers and a sharp soliton at the center $(8,8)$. The top row shows band structures, the bottom row shows the corresponding density of states $g(E)$ for different values of the magnitude of the alternating onsite energies as (a), (f) $U/\gamma_{1}=0.0$, (b), (g) $U/\gamma_{1}=0.25$, (c), (h) $U/\gamma_{1}=0.5$, (d), (i) $U/\gamma_{1}=1.0$, and (e), (j) $U/\gamma_{1}=2.0$. In the band structure plots, horizontal arrows show the energies at $k_{x}=0$ of the states localized on the soliton. In all plots, parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [35], $a=2.46$Å [36]. For the band structures, $k_{y}=0$, and, for the density of states, $\delta=0.01\gamma_{1}$.
V.2 ${}^{\prime}$
@@ V.2.1 Faultless system
For bulk, faultless ${}^{\prime}$ stacking, using the approximation for the intralayer hopping (23) with $k<k_{c}$, we can use a unit cell of two atoms, either $A_{i}$, $B_{i}$ on an odd layer or $B_{i}$, $A_{i}$ on an even layer, to give the Hamiltonian
@@ V.2.2 Single soliton
For small values of $U/\gamma_{1}$, there are crossings between levels, including the two low-energy levels, Fig. 11(b), whereby the energies appear to oscillate as a function of $k$. Using spatial inversion symmetry, it is possible to block diagonalize the Hamiltonian (6) into a block with even parity states and a block with odd parity states. The odd parity block has eigenstates which are the electron-hole reflection ($E\rightarrow-E$) of the even parity eigenstates. Each individual block describes a system similar to ${}^{\prime}$ stacking with a spectrum having anticrossings at $U\neq 0$ and a single soliton level. For small $U/\gamma_{1}$, the soliton level crosses zero energy, at which point its electron-hole partner (from the opposite parity block) crosses zero energy in the opposite direction. Oscillations are caused by anticrossings with levels from the same block. The anticrossings, and resulting oscillations, are reduced as $U/\gamma_{1}$ increases because the other levels (in the bulk conduction and valence bands) move further away in energy and the soliton-antisoliton pair move away from zero energy and separate. This picture holds in the presence of spatial inversion symmetry. When the soliton is off-center, the band structure is similar, but the level crossings between the levels with even or odd parity are replaced by anticrossings.
@@ VI Smooth solitons
Figure 12: Band structures and density of states $g(E)$ for smooth solitons with width $\zeta=8$ at the center $(8,8)$ of a system with $N=16$ layers for ABA, ABC, and ${}^{\prime}$ stacking. In the band structure plots, horizontal arrows show the energies at $k_{x}=0$ of the states localized on the soliton. There is a different axis scale for ${}^{\prime}$ stacking. In all plots, parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [35], $a=2.46$Å [36]. For the band structures, $k_{y}=0$, and, for the density of states, $\delta=0.01\gamma_{1}$.
Atomically sharp solitons support localized states within the bulk band gap for ABC, ABA and ${}^{\prime}$ stacking, Fig. 1. Now we will describe the properties of smooth solitons with a finite width for these three stacking types. Band structures and density of states are determined as described in the methodology, Section II, by numerically diagonalizing Hamiltonian (6) after replacing the diagonal elements with a smooth soliton texture. Specifically, for ABC and ABA stacking, we model the onsite energy of an A site on the $j$th layer, where $j=1,2,\ldots,N$, as
@@
where $U$ is the magnitude of the texture at infinity, $m$ is the number of layers below the fault, and $\zeta$ is the soliton width in dimensionless units, i.e., measured in units of the interlayer spacing. For ${}^{\prime}$ stacking,
@@ VI.2 ABA stacking
VI.3 ${}^{\prime}$ stacking
For ${}^{\prime}$ stacking, Fig. 12 shows that there are a pair of bands (soliton and antisoliton) near zero energy, as for a sharp soliton, but the major difference is that these two bands do not separate as $U/\gamma_{1}$ increases, but remain near zero energy as flat bands with a large extent up to $k\approx 2k_{c}$.
@@ VII Single layer defects
Figure 14: Band structures and density of states $g(E)$ for single layer defects near the center $(7,1,8)$ of a system with $N=16$ layers for ABA, ABC, and ${}^{\prime}$ stacking. In the band structure plots, horizontal arrows show the energies at $k_{x}=0$ of the states localized on the defect. There is a different axis scale for ${}^{\prime}$ stacking, and the other stackings at large values of $U/\gamma_{1}$. In all plots, parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [35], $a=2.46$Å [36]. For the band structures, $k_{y}=0$, and, for the density of states, $\delta=0.01\gamma_{1}$.
So far, we have considered solitons which consist of a change in the texture of onsite energies without changing the interatomic hopping, and they may be sharp on the atomic scale or smooth with a nonzero width $\zeta$. There are many other types of stacking faults one could consider. An example, which also involves onsite energies, is to reverse the signs of the onsite energies on a single layer only. This doesn’t change the texture of the onsite energies either side of the defect, and we refer to this as a ‘single layer defect’, not a soliton. Such defects were modeled for AA, ${}^{\prime}$, ABA and ${}^{\prime}$ stacking using density functional theory (DFT) in the context of hexagonal boron nitride [33], and we will only discuss them briefly here, for ABC, ABA and ${}^{\prime}$ stacking. The lattice structures, with the defects, are shown schematically in the top panels of Fig. 14 for a system we denote as $(3,1,4)$ indicating a system of $N=8$ layers consisting of $3$ layers (at the bottom) with regular stacking, followed by a single layer defect, followed by $4$ layers (at the top) with regular stacking. Band structures and density of states for ABC, ABA, and ${}^{\prime}$ stacking, are shown in the remaining panels in Fig. 14 for a system with a larger number of layers $N=16$ and a defect near its center $(7,1,8)$.
@@ VII.1 ABC stacking
The behavior at $k=0$ may be understand because the system separates into dimers and isolated atoms there. There are the two isolated surface states with energies $E=\pm U$, $N-3$ dimers each with energies $E=\pm\sqrt{U^{2}+\gamma_{1}^{2}}$, and, near the defect, a dimer with onsite energies $-U$ giving $E=-U\pm\gamma_{1}$ and a dimer with onsite energies $+U$ giving $E=+U\pm\gamma_{1}$. These four energies are shown with the horizontal arrows in Fig. 14.
@@ VII.2 ABA stacking
For ABA stacking, Fig. 14 shows that there is a single defect state within the bulk band gap. For small $U/\gamma_{1}$, it is near the bottom of the conduction band for $k=0$, crossing the band gap to move into the valence band at $k\gg k_{c}$, whereas, for larger $U/\gamma_{1}$, it is always at negative energy. The lattice structure, Fig. 14(b), shows there are three consecutive sites in the $N$-mer with the same onsite energies $E=-U$. For $k=0$ and $U\gg\gamma_{1}$, this trimer gives energies $E=-U$ and $E=-U\pm\sqrt{2}\gamma_{1}$ (as indicated by the horizontal arrows in Fig. 14). The level $E=-U+\sqrt{2}\gamma_{1}$ is the one within the band gap.
VII.3 ${}^{\prime}$ stacking
For ${}^{\prime}$ stacking, Fig. 14 shows that there are generally two defect states within the band gap, one merges with the conduction band for $k\gg k_{c}$, the other merges with the valence band. The lattice structure, Fig. 14, shows the ladder structure consisting of two coupled $N$-mers. In one of them, there is a trimer of states with onsite energy $E=+U$ near the defect, and, in the other, there is a trimer of states with onsite energy $E=-U$ near the defect. Together, they contribute six states which are, at $k=0$ and $U\gg\gamma_{1}$, $E=\pm U$, $E=-U\pm\sqrt{2}\gamma_{1}$, and $E=U\pm\sqrt{2}\gamma_{1}$ (as indicated by the horizontal arrows in Fig. 14). The two levels $E=-U+\sqrt{2}\gamma_{1}$ and $E=U-\sqrt{2}\gamma_{1}$ are within the band gap.
@@ VIII The role of interlayer disorder
We have shown that solitons, both atomically-sharp and smooth in position space, are able to support localized states with energies within the bulk band gap for ABC, ABA and ${}^{\prime}$ stacking. The soliton bands are generally dispersive as a function of the in-plane wave vector ${\bf k}$ so that they give rise to a non-zero density of states without any particularly sharp features. The exception are smooth solitons for ${}^{\prime}$ stacking which give flat bands yielding a narrow peak at zero energy in the density of states, Fig. 12(x).
For certain values of the in-plane wave vector ${\bf k}$, the lattice structure of these systems may be related by dimensional reduction to that of the CDW model [41, 42, 15, 43, 44, 16] which has nonsymmorphic chiral symmetry. The influence of symmetry-breaking disorder has been considered previously for the CDW model [44] so here we consider it only for the most interesting case of ${}^{\prime}$ stacking, for both atomically-sharp and smooth solitons.
To focus on the topology related to the CDW model, we consider random tight-binding parameters in the perpendicular-to-layer direction while preserving translational invariance in the in-plane direction. We consider four types of disorder: (i) “Sharp onsite disorder” (diagonal) which gives an additional contribution $\delta\epsilon_{A,j}$ to the onsite energy of an A site on the $j$th layer, where $j=1,2,\ldots,N$, which is drawn randomly from a uniform distribution $-W\leq\delta\epsilon_{A,j}\leq W$ with disorder strength $W$. We maintain charge neutrality within each layer so that $\delta\epsilon_{B,j}=-\delta\epsilon_{A,j}$ for all layers $j$. (ii) “Sharp hopping disorder” (off-diagonal) where the interlayer coupling takes random values $\gamma_{1}+\delta_{n}$ where $n=1,2,\ldots,N-1$ indexes the $N-1$ interlayer spaces and each $\delta_{n}$ takes a value drawn randomly from a uniform distribution $-W\leq\delta_{n}\leq W$. For ${}^{\prime}$ stacking, the values of coupling between A atoms and between B atoms for any pair of adjacent layers are identical, so that there are only $N-1$ random values of $\delta_{n}$ in total. (iii) “Smooth onsite disorder” (diagonal) with an additional contribution $\delta\epsilon_{A,j}$ to the onsite energy of an A site on the $j$th layer described by a Gaussian-correlated potential [50, 51] as given by
@@
Figure 15: Dependence of the disorder-averaged density of states $\langle g(E)\rangle$ on energy $E$ for an atomically-sharp soliton at the center $(8,8)$ of a system with ${}^{\prime}$ stacking and $N=16$ layers. For all plots, the onsite energy is $U=2.0\gamma_{1}$, the disorder strength is $W=0.5\gamma_{1}$, and $\langle g(E)\rangle$ is determined using Eq. (20) with broadening $\delta=0.01\gamma_{1}$ and $20$ disorder realizations. Other parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [35], $a=2.46$Å [36]. (a) is for atomically-sharp onsite disorder (black), (b) is for sharp hopping disorder (magenta), (c) is for Gaussian-correlated onsite disorder (122) with correlation length $\eta=4$ (red), and (d) is for Gaussian-correlated hopping disorder with $\eta=4$ (blue).
@@
Figure 16: Dependence of the disorder-averaged density of states $\langle g(E)\rangle$ on energy $E$ for a smooth soliton of width $\zeta=8$ at the center $(8,8)$ of a system with ${}^{\prime}$ stacking and $N=16$ layers. For all plots, the onsite energy is $U=2.0\gamma_{1}$, the disorder strength is $W=0.5\gamma_{1}$, and $\langle g(E)\rangle$ is determined using Eq. (20) with broadening $\delta=0.01\gamma_{1}$ and $20$ disorder realizations. Other parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [35], $a=2.46$Å [36]. (a) is for atomically-sharp onsite disorder (black), (b) is for sharp hopping disorder (magenta), (c) is for Gaussian-correlated onsite disorder (122) with correlation length $\eta=4$ (red), and (d) is for Gaussian-correlated hopping disorder with $\eta=4$ (blue).
@@
The normalisation with the square root factor in Eq. (122) is used so that smooth disorder interpolates between sharp disorder for $\eta\ll 1$ and sample-to-sample parameter variations [44] for $\eta\gg N$. The latter are parameter values that are spatially uniform across a single member of the ensemble, i.e., $\delta\epsilon_{A,j}=\delta\epsilon_{A}$ for all $j$, but that differ from sample to sample with $\delta\epsilon_{A}$ drawn randomly from a uniform distribution $-W\leq\delta\epsilon_{A}\leq W$. Smooth solitons in the CDW model are fairly robust to parameter variations [44] because the nonsymmorphic chiral symmetry holds in the continuum limit (of a smooth soliton with parameter variations). However, the equivalence of ${}^{\prime}$ stacking with the CDW model only holds at $k=0$, so smooth solitons for ${}^{\prime}$ stacking are not expected to be as robust. As parameter variations were studied in detail for the CDW model [44], we consider instead the case of a smooth potential with a finite range $1\lesssim\eta\lesssim N$.
@@
For ${}^{\prime}$ stacking and an atomically sharp soliton in the absence of disorder, Fig 11(j), the density of states has a step-like nonzero density of states within the bulk band gap $-U\leq E\leq U$ due to the soliton bands, Fig 11(e). In addition, there are also sharp peaks in the density of states at the bulk band edges, $E=\pm U$. Figure 15 shows the disorder-averaged density of states $\langle g(E)\rangle$ for an atomically-sharp soliton at the center $(8,8)$ of a system with ${}^{\prime}$ stacking and $N=16$ layers. We consider the example of onsite energy $U=2.0\gamma_{1}$ and disorder strength $W=0.5\gamma_{1}$. Figure 15 shows that all types of disorder tend to smooth out the step-like feature due to the soliton states within the bulk band gap. The sharp peaks at $E=\pm U$ at the band edges are quite robust to disorder, as they are only smoothed out by atomically-sharp onsite disorder, Figure 15(a).

### v2

@@ IV.1.3 Soliton-antisoliton pair
Figure 9: Low-energy band structures for ABA stacking. Each column shows a different value of the onsite energies $U$. The top row shows a faultless system with $N=16$ layers, whereas the second row shows $N=16$ layers with a sharp fault at its center $(8,8)$. In all plots, parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [40], $a=2.46$Å [41].
IV.2 ABA
@@ IV.2.1 Faultless system
For bulk, faultless ABA stacking, using the approximation for the intralayer hopping (8) with $k<k_{c}$, we can use a unit cell of two atoms, either $A_{i}$, $B_{i}$ on an odd layer or $B_{i}$, $A_{i}$ on an even layer, to give the Hamiltonian
@@ IV.2.2 Single soliton
Band structures for ABA stacking with a sharp fault at its center are shown in Fig. 9 (bottom row). At $k=0$, the system separates into isolated atoms with energies ±$\pm U$ plus the $N$-mer chain. Now the $N$-mer chain contains a fault, so it consists of a monoatomic chain with onsite energies $-U$ and nearest neighbor hopping $\gamma_{1}$ connected to a monoatomic chain with onsite energies $+U$ and nearest neighbor hopping $\gamma_{1}$. The former has a bulk band $E(q_{z})=-U+2\gamma_{1}\cos(q_{z}d)$, the latter has a bulk band $E(q_{z})=U+2\gamma_{1}\cos(q_{z}d)$, i.e., the system consists of one part with a band centered on energy $+U$ and of bandwidth $4\gamma_{1}$, connected to another part with a band centered on energy $-U$ and of bandwidth $4\gamma_{1}$. For $U\leq 2\gamma_{1}$, the two bands overlap leading, in a system of a finite number of layers, to overlapping bands with anticrossings at low energy as in Fig. 9(h) to Fig. 9(j). For $U>2\gamma_{1}$, the two bands do not overlap, giving an overall band gap as in Fig. 9(l).
Although the presence of a fault for ABA stacking creates changes in the band structure, Fig. 9, there are no states localized on the fault. This can be understood by considering the system at $k=0$: The fault divides two sections of monoatomic chains with different onsite energies ($+U$ or $-U$) and, thus, the two sections are not degenerate, nor do they have a Dirac-like low-energy continuum description [25].
@@ V.1.2 Single soliton
As with ABA stacking, although the presence of a fault for AA stacking creates changes in the band structure, Fig. 10, there are no states localized on the fault. The term describing hopping along the ladder, $2\gamma_{1}\cos(q_{z}d)$, appears on the main diagonal in the ladder Hamiltonian (36) and it breaks chiral symmetry. The trivial nature of the fault may be further understood by considering the limit $k=0$ when the system separates into monoatomic chains which do not have a Dirac-like low-energy continuum description [25].
Figure 11: AA stacking with $N=16$ layers and a sharp soliton at the center $(8,8)$. The top row shows band structures, the bottom row shows the corresponding density of states $g(E)$ for different values of the magnitude of the alternating onsite energies as (a), (f) $U/\gamma_{1}=0.0$, (b), (g) $U/\gamma_{1}=0.25$, (c), (h) $U/\gamma_{1}=0.5$, (d), (i) $U/\gamma_{1}=1.0$, and (e), (j) $U/\gamma_{1}=2.0$. In the band structure plots, horizontal arrows show the energies at $k_{x}=0$ of the states localized on the soliton. In all plots, parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [40], $a=2.46$Å [41]. For the band structures, $k_{y}=0$, and, for the density of states, $\delta=0.01\gamma_{1}$.
V.2 AA
@@ V.2.1 Faultless system
For bulk, faultless AA stacking, using the approximation for the intralayer hopping (8) with $k<k_{c}$, we can use a unit cell of two atoms, either $A_{i}$, $B_{i}$ on an odd layer or $B_{i}$, $A_{i}$ on an even layer, to give the Hamiltonian
@@ V.2.2 Single soliton
For small values of $U/\gamma_{1}$, there are crossings between levels, including the two low-energy levels, Fig. 11(b), whereby the energies appear to oscillate as a function of $k$. Using spatial inversion symmetry, it is possible to block diagonalize the Hamiltonian (1) into a block with even parity states and a block with odd parity states. The odd parity block has eigenstates which are the electron-hole reflection ($E\rightarrow-E$) of the even parity eigenstates. Each individual block describes a system similar to AA stacking with a spectrum having anticrossings at $U\neq 0$ and a single soliton level. For small $U/\gamma_{1}$, the soliton level crosses zero energy, at which point its electron-hole partner (from the opposite parity block) crosses zero energy in the opposite direction. Oscillations are caused by anticrossings with levels from the same block. The anticrossings, and resulting oscillations, are reduced as $U/\gamma_{1}$ increases because the other levels (in the bulk conduction and valence bands) move further away in energy and the soliton-antisoliton pair move away from zero energy and separate. This picture holds in the presence of spatial inversion symmetry. When the soliton is off-center, the band structure is similar, but the level crossings between the levels with even or odd parity are replaced by anticrossings.
@@ VI Smooth solitons
Figure 12: Band structures and density of states $g(E)$ for smooth solitons with width $\zeta=8$ at the center $(8,8)$ of a system with $N=16$ layers for ABA, ABC, and AA stacking. In the band structure plots, horizontal arrows show the energies at $k_{x}=0$ of the states localized on the soliton. There is a different axis scale for AA stacking. In all plots, parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [40], $a=2.46$Å [41]. For the band structures, $k_{y}=0$, and, for the density of states, $\delta=0.01\gamma_{1}$.
Atomically sharp solitons support localized states within the bulk band gap for ABC, ABA and AA stacking, Fig. 1. Now we will describe the properties of smooth solitons with a finite width for these three stacking types. Band structures and density of states are determined as described in the methodology, Section II, by numerically diagonalizing Hamiltonian (1) after replacing the diagonal elements with a smooth soliton texture. Specifically, for ABC and ABA stacking, we model the onsite energy of an A site on the $j$th layer, where $j=1,2,\ldots,N$, as
@@
where $U$ is the magnitude of the texture at infinity, $m$ is the number of layers below the fault, and $\zeta$ is the soliton width in dimensionless units, i.e., measured in units of the interlayer spacing. For AA stacking,
@@ VI.2 ABA stacking
VI.3 AA stacking
For AA stacking, Fig. 12 shows that there are a pair of bands (soliton and antisoliton) near zero energy, as for a sharp soliton, but the major difference is that these two bands do not separate as $U/\gamma_{1}$ increases, but remain near zero energy as flat bands with a large extent up to $k\approx 2k_{c}$.
@@ VII Single layer defects
Figure 14: Band structures and density of states $g(E)$ for single layer defects near the center $(7,1,8)$ of a system with $N=16$ layers for ABA, ABC, and AA stacking. In the band structure plots, horizontal arrows show the energies at $k_{x}=0$ of the states localized on the defect. There is a different axis scale for AA stacking, and the other stackings at large values of $U/\gamma_{1}$. In all plots, parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [40], $a=2.46$Å [41]. For the band structures, $k_{y}=0$, and, for the density of states, $\delta=0.01\gamma_{1}$.
So far, we have considered solitons which consist of a change in the texture of onsite energies without changing the interatomic hopping, and they may be sharp on the atomic scale or smooth with a nonzero width $\zeta$. There are many other types of stacking faults one could consider. An example, which also involves onsite energies, is to reverse the signs of the onsite energies on a single layer only. This doesn’t change the texture of the onsite energies either side of the defect, and we refer to this as a ‘single layer defect’, not a soliton. Such defects were modeled for AA, AA, ABA and ABA stacking using density functional theory (DFT) in the context of hexagonal boron nitride [38], and we will only discuss them briefly here, for ABC, ABA and AA stacking. The lattice structures, with the defects, are shown schematically in the top panels of Fig. 14 for a system we denote as $(3,1,4)$ indicating a system of $N=8$ layers consisting of $3$ layers (at the bottom) with regular stacking, followed by a single layer defect, followed by $4$ layers (at the top) with regular stacking. Band structures and density of states for ABC, ABA, and AA stacking, are shown in the remaining panels in Fig. 14 for a system with a larger number of layers $N=16$ and a defect near its center $(7,1,8)$.
@@ VII.1 ABC stacking
The behavior at $k=0$ may be understand because the system separates into dimers and isolated atoms there. There are the two isolated surface states with energies $E=\pm U$, $N-3$ dimers each with energies $E=\pm\sqrt{U^{2}+\gamma_{1}^{2}}$, and, near the defect, a dimer with onsite energies $-U$ giving $E=-U\pm\gamma_{1}$ and a dimer with onsite energies $+U$ giving $E=+U\pm\gamma_{1}$. These four energies are shown with the horizontal arrows in Fig. 14. The presence of two defect states within the band gap is in stark contrast to a single soliton which only supports one state within the gap, Fig. 2. Instead, a single layer defect has a band structure very similar to that of a soliton-antisoliton pair, Fig. 5, because the latter can also be viewed as a defect, albeit of larger spatial extent than just one layer, which doesn’t change the texture of the onsite energies either side of the defect.
@@ VII.2 ABA stacking
For ABA stacking, Fig. 14 shows that there is a single defect state within the bulk band gap. For small $U/\gamma_{1}$, it is near the bottom of the conduction band for $k=0$, crossing the band gap to move into the valence band at $k\gg k_{c}$, whereas, for larger $U/\gamma_{1}$, it is always at negative energy. The lattice structure, Fig. 14(b), shows there are three consecutive sites in the $N$-mer with the same onsite energies $E=-U$. For $k=0$ and $U\gg\gamma_{1}$, this trimer gives energies $E=-U$ and $E=-U\pm\sqrt{2}\gamma_{1}$ (as indicated by the horizontal arrows in Fig. 14). The level $E=-U+\sqrt{2}\gamma_{1}$ is the one within the band gap. A single soliton also creates a single level within the band gap, Fig. 7, and the band structure and density of states of a single soliton and a defect are similar, particularly for $U\gg\gamma_{1}$. For $U\ll\gamma_{1}$, the defect state extends across most of the band gap, creating a wide energy region with a nonzero density of states, unlike the soliton state.
VII.3 AA stacking
For AA stacking, Fig. 14 shows that there are generally two defect states within the band gap, one merges with the conduction band for $k\gg k_{c}$, the other merges with the valence band. The lattice structure, Fig. 14, shows the ladder structure consisting of two coupled $N$-mers. In one of them, there is a trimer of states with onsite energy $E=+U$ near the defect, and, in the other, there is a trimer of states with onsite energy $E=-U$ near the defect. Together, they contribute six states which are, at $k=0$ and $U\gg\gamma_{1}$, $E=\pm U$, $E=-U\pm\sqrt{2}\gamma_{1}$, and $E=U\pm\sqrt{2}\gamma_{1}$ (as indicated by the horizontal arrows in Fig. 14). The two levels $E=-U+\sqrt{2}\gamma_{1}$ and $E=U-\sqrt{2}\gamma_{1}$ are within the band gap. A single soliton also creates two levels within the band gap, Fig. 11, and the band structure and density of states of a single soliton and a defect are similar. There are differences such as the hybridization of the two levels with the levels being close together and strongly hybridized for $U\ll\gamma_{1}$ in the soliton and for $U\approx\gamma_{1}$ in the defect.
@@ VIII The role of interlayer disorder
We have shown that solitons, both atomically-sharp and smooth in position space, are able to support localized states with energies within the bulk band gap for ABC, ABA and AA stacking. The soliton bands are generally dispersive as a function of the in-plane wave vector ${\bf k}$ so that they give rise to a non-zero density of states without any particularly sharp features. The exception are smooth solitons for AA stacking which give flat bands yielding a narrow peak at zero energy in the density of states, Fig. 12(x).
For certain values of the in-plane wave vector ${\bf k}$, the lattice structure of these systems may be related by dimensional reduction to that of the CDW model [33, 34, 15, 35, 36, 16] which has nonsymmorphic chiral symmetry. The influence of symmetry-breaking disorder has been considered previously for the CDW model [36] so here we consider it only for the most interesting case of AA stacking, for both atomically-sharp and smooth solitons.
To focus on the topology related to the CDW model, we consider random tight-binding parameters in the perpendicular-to-layer direction while preserving translational invariance in the in-plane direction. We consider four types of disorder: (i) “Sharp onsite disorder” (diagonal) which gives an additional contribution $\delta\epsilon_{A,j}$ to the onsite energy of an A site on the $j$th layer, where $j=1,2,\ldots,N$, which is drawn randomly from a uniform distribution $-W\leq\delta\epsilon_{A,j}\leq W$ with disorder strength $W$. We maintain charge neutrality within each layer so that $\delta\epsilon_{B,j}=-\delta\epsilon_{A,j}$ for all layers $j$. (ii) “Sharp hopping disorder” (off-diagonal) where the interlayer coupling takes random values $\gamma_{1}+\delta_{n}$ where $n=1,2,\ldots,N-1$ indexes the $N-1$ interlayer spaces and each $\delta_{n}$ takes a value drawn randomly from a uniform distribution $-W\leq\delta_{n}\leq W$. For AA stacking, the values of coupling between A atoms and between B atoms for any pair of adjacent layers are identical, so that there are only $N-1$ random values of $\delta_{n}$ in total. (iii) “Smooth onsite disorder” (diagonal) with an additional contribution $\delta\epsilon_{A,j}$ to the onsite energy of an A site on the $j$th layer described by a Gaussian-correlated potential [49, 50] as given by
@@
Figure 15: Dependence of the disorder-averaged density of states $\langle g(E)\rangle$ on energy $E$ for an atomically-sharp soliton at the center $(8,8)$ of a system with AA stacking and $N=16$ layers. For all plots, the onsite energy is $U=2.0\gamma_{1}$, the disorder strength is $W=0.5\gamma_{1}$, and $\langle g(E)\rangle$ is determined using Eq. (7) with broadening $\delta=0.01\gamma_{1}$ and $20$ disorder realizations. Other parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [40], $a=2.46$Å [41]. (a) is for atomically-sharp onsite disorder (black), (b) is for sharp hopping disorder (magenta), (c) is for Gaussian-correlated onsite disorder (42) with correlation length $\eta=4$ (red), and (d) is for Gaussian-correlated hopping disorder with $\eta=4$ (blue).
@@
Figure 16: Dependence of the disorder-averaged density of states $\langle g(E)\rangle$ on energy $E$ for a smooth soliton of width $\zeta=8$ at the center $(8,8)$ of a system with AA stacking and $N=16$ layers. For all plots, the onsite energy is $U=2.0\gamma_{1}$, the disorder strength is $W=0.5\gamma_{1}$, and $\langle g(E)\rangle$ is determined using Eq. (7) with broadening $\delta=0.01\gamma_{1}$ and $20$ disorder realizations. Other parameter values are $\gamma_{0}=3.16\,$eV, $\gamma_{1}=0.381\,$eV [40], $a=2.46$Å [41]. (a) is for atomically-sharp onsite disorder (black), (b) is for sharp hopping disorder (magenta), (c) is for Gaussian-correlated onsite disorder (42) with correlation length $\eta=4$ (red), and (d) is for Gaussian-correlated hopping disorder with $\eta=4$ (blue).
@@
The normalisation with the square root factor in Eq. (42) is used so that smooth disorder interpolates between sharp disorder for $\eta\ll 1$ and sample-to-sample parameter variations [36] for $\eta\gg N$. The latter are parameter values that are spatially uniform across a single member of the ensemble, i.e., $\delta\epsilon_{A,j}=\delta\epsilon_{A}$ for all $j$, but that differ from sample to sample with $\delta\epsilon_{A}$ drawn randomly from a uniform distribution $-W\leq\delta\epsilon_{A}\leq W$. Smooth solitons in the CDW model are fairly robust to parameter variations [36] because the nonsymmorphic chiral symmetry holds in the continuum limit (of a smooth soliton with parameter variations). However, the equivalence of AA stacking with the CDW model only holds at $k=0$, so smooth solitons for AA stacking are not expected to be as robust. As parameter variations were studied in detail for the CDW model [36], we consider instead the case of a smooth potential with a finite range $1\lesssim\eta\lesssim N$.
@@
For AA stacking and an atomically sharp soliton in the absence of disorder, Fig 11(j), the density of states has a step-like nonzero density of states within the bulk band gap $-U\leq E\leq U$ due to the soliton bands, Fig 11(e). In addition, there are also sharp peaks in the density of states at the bulk band edges, $E=\pm U$. Figure 15 shows the disorder-averaged density of states $\langle g(E)\rangle$ for an atomically-sharp soliton at the center $(8,8)$ of a system with AA stacking and $N=16$ layers. We consider the example of onsite energy $U=2.0\gamma_{1}$ and disorder strength $W=0.5\gamma_{1}$. Figure 15 shows that all types of disorder tend to smooth out the step-like feature due to the soliton states within the bulk band gap. The sharp peaks at $E=\pm U$ at the band edges are quite robust to disorder, as they are only smoothed out by atomically-sharp onsite disorder, Figure 15(a).